# Centre du Parc, Rue Marconi 19, CH-1920 Martigny, Switzerland

### v1

@@ A.4 Experiments
Obstacle Avoidance Push. Cyl. Push. Box Push. Must. Sphere man. Sinus man.
$H$ $15$ $32$ $20$ $20$ $50$ $30$
$\bm{\Sigma}_{\text{task}}$ $0.125\bm{I}$ $0.3\bm{I}$ $0.05\bm{I}$ $0.05\bm{I}$ $0.1\bm{I}$ $0.05\bm{I}$
${\bm{\Sigma}_{z}}$ $10\bm{I}$ $0.5\bm{I}$ $0.01\bm{I}$ $0.05\bm{I}$ $0.2\bm{I}$ $10\bm{I}$
$\beta$ $0.05$ $0.05$ $0.05$ $0.05$ $0.05$ $0.05$
$dt[s]$ $0.1$ $0.05$ $0.05$ $0.05$ $0.1$ $0.1$
$u_{\max}[\frac{m}{s}]$ $1$ $1$ $1$ $1$ $1$ $1$
$x_{\max}[m]$ $1.25$ $1$ $1$ $0.4$ $0.25$ $0.5$
$T_{{\text{succ.}}}$ $100$ $500$ $500$ $400$ $400$ $200$
$c_{\text{succ.}}$ $10^{30}$ $10^{4}$ $10^{4}$ $10^{3}$ $10^{6}$ $010^{3}$ 0
@@
where the notation $\hat{\cdot}$ denotes the values of the cost functions before normalization. Additionally, we incorporate an extra step by assuming one of the samples to have zero action. This facilitates the system to halt in case all samples become infeasible. Further explanation regarding the employed cost functions in these experiments is provided in the following.
A.4.1 Obstacle avoidance:
@@
A.4.2 Planar-pushing:
We have modeled the robot end-effector as a 2D point mass under the control of a velocity controller. For each object, we have utilized distinct parameters, detailed in Table LABEL:tab:exp_data, and is subjected to unique cost functions, elaborated upon subsequently.
@@ A.4.2 Planar-pushing:
[…]
Pushing a Mustard Bottle and a Box We use the same cost function for these two tasks which aimed at reaching a target position $\bm{p}_{\text{target}}$ and desired yaw angle $o^{\text{d}}$ as
@@
A.4.3 Moving on Manifolds:
In this task, our objective is to control the robot end-effector’s position within a predefined area. In the sphere manifold scenario, the area is bounded by two spheres: the inner sphere has a radius of 15 cm, while the outer sphere has a radius of 20 cm. For the sinusoidal wave, the area is defined as the region inside $\{z\mid z\in 0.1\sin(4y\pi)\pm 0.03,\quad y\in[-1,1]\}$. During both the learning and planning phases, collision detection incorporates a 1-cm margin to ensure a safety buffer.
@@ A.4.3 Moving on Manifolds:
A.5 Real-robot Experiment

### v2

@@ A.7 Experiments
Obstacle Avoidance Push. Cyl. Push. Box Push. Must. Sphere man. Sinus man. Whole-body
$H$ $15$ $32$ 2020 2020 $50$ $30$ $15$
$\bm{\Sigma}_{\text{task}}$ $0.125\bm{I}$ $3\bm{I} 0.0$$5\bm{I} 0.0$$05\bm{I}$ $1\bm{I} 0.0$$5\bm{I} 0.0$$05\bm{I}$
$\bm{\Sigma}_{z}$} $10\bm{I}$ $5\bm{I} 0.0$$1\bm{I} 0.0$$05\bm{I}$ $0.2\bm{I}$ $10\bm{I}$ -
$\beta$ $0.05$ $0.05$ $0.05$ $0.05$ $0.05$ $0.05$ $0.01$
$dt[s]$ $0.1$ $0.05$ $0.05$ $0.05$ $0.1$ $0.1$ $0.1$
$u_{\max}[\frac{m}{s}]$ 11 11 11 11 11 11 Real Limits
$x_{\max}[m]$ $1.25$ 11 11 $0.4$ $0.25$ $0.5$ Real Limits
$T_{{\text{succ.}}}$ $100$ $500$ $500$ $400$ $400$ $200$ $25$
$c_{\text{succ.}}$ $10^{30}$ $10^{4}$ $10^{4}$ $10^{3}$ $3\times 10^{5}$ $10^{30}$ $10^{6}$
$N_{\text{diffuse}}$ - - - - - - 33
$\text{f}_{\text{denoise}}$ - - - - - - $0.22$
$\text{f}_{\text{horizon}}$ - - - - - - $0.95$
@@
where the notation $\hat{(\cdot)}$ denotes the values of the cost functions before normalization. Additionally, we incorporate an extra step by assuming one of the samples to have zero action. This facilitates the system to halt in case all samples become infeasible. Further explanation regarding the employed cost functions in these experiments is provided in the following.
A.7.1 Obstacle avoidance:
@@
For Expert 2, we have utilized the following feasibility distribution
$p_{\text{feas}}(\hat{\bm{x}}_{t},\hat{\bm{u}}_{t})\;=\;\bm{1}\!\bigl\{\hat{\bm{x}}_{t}\in\mathcal{X},\;\hat{\bm{x}}_{t+1}\in\mathcal{X},\;(\hat{\bm{x}}_{t},\hat{\bm{x}}_{t+1})\cap\mathcal{O}=\emptyset\bigr\},$
where, $\mathcal{X}$ denotes the allowable state‐space (inside the boundaries), $\mathcal{O}$ the obstacle set, and $\bm{1}\{\cdot\}$ is the indicator function.
@@
A.7.2 Planar-pushing:
We have modeled the robot end-effector as a 2D point mass under the control of a velocity controller. For each object, we have utilized distinct parameters, detailed in Table LABEL:tab:exp_data, and is subjected to unique cost functions, elaborated upon subsequently. For Expert 2, however, all objects share the same feasibility distribution
$p_{\text{feas}}(\hat{\bm{x}}_{t},\hat{\bm{u}}_{t})=\bm{1}\Bigl\{\hat{\bm{x}}^{\text{obj.}}\in\mathcal{X}\land\Big(\Delta_{o}>0\;\lor\;{\Delta_{d}}>0\Big)\Bigr\},$
where
$\begin{gathered}\Delta_{o}=\|\hat{\bm{x}}^{\text{obj.}}_{t+1}-\hat{\bm{x}}^{\text{obj.}}_{t}\|,\\ \Delta_{d}=\|\hat{\bm{x}}^{\text{rob.}}_{t}-\hat{\bm{x}}^{\text{obj.}}_{t}\|-\|\hat{\bm{x}}^{\text{rob.}}_{t+1}-\hat{\bm{x}}^{\text{obj.}}_{t+1}\|.\end{gathered}$
Here, $\hat{\bm{x}}^{\mathrm{obj}}$ and $\hat{\bm{x}}^{\mathrm{rob}}$ denote the object and robot states, respectively. This distribution assigns high probability only to samples in which the object remains within the feasible region and either (i) the object moves ($\Delta_{o}>0$) or (ii) the robot–object distance decreases ($\Delta_{d}>0$).
@@ A.7.2 Planar-pushing:
where functions $\text{coll\_o}(\cdot)$ and $\text{coll\_b}(\cdot)$ determine collision with the obstacles and the boundary, respectively, returning 1 in case of collision and 0 otherwise. Additionally, $\text{SDF}(\cdot)$ calculates the signed distance value between the robot and the agent. Similarly to the obstacle avoidance experiment, $r_{t+h}$ serves as an indicator to verify whether the agent reached the target before the time step $t+h$. We define the target as reached when the distance between the agent and the target point is less than 5 centimeters.
Pushing a Mustard Bottle and a Box: We use the same cost function for these two tasks which aimed at reaching a target position $\bm{p}_{\text{target}}$ and desired yaw angle $o^{\text{d}}$ as
@@
A.7.3 Tube Tracking:
In this task, our objective is to control the robot end-effector’s position within a predefined area. In the sphere manifold scenario, the area is bounded by two spheres: the inner sphere has a radius of 15 cm, while the outer sphere has a radius of 20 cm. For the sinusoidal wave, the area is defined as the region inside $\{(x,y,z)\mid z\in 0.1\sin(5y\pi)\pm 0.03,\quad x\in[0.2,1.0],\quad y\in[-0.5,0.5]\}$. During both the learning and planning phases, collision detection incorporates a 1-cm margin to ensure a safety buffer.
@@ A.7.3 Tube Tracking:
For Expert 2, we employ the following feasibility distribution
$p_{\text{feas}}(\hat{\bm{x}}_{t},\hat{\bm{u}}_{t})=\bm{1}\!\Bigl\{\hat{\bm{x}}_{t}\in\mathcal{X},\;\hat{\bm{x}}_{t+1}\in\mathcal{X},\;\hat{\bm{x}}_{t},\hat{\bm{x}}_{t+1}\in\mathcal{M}_{\epsilon}\Bigr\},$
where
$\mathcal{M}_{\epsilon}=\{\,\bm{x}\mid\mathrm{dist}(\bm{x},\mathcal{M})\leq\epsilon\}$
is an $\epsilon$-margin around the manifold $\mathcal{M}$. This ensures that both the current and next states lie within the admissible state space $\mathcal{X}$ and remain close to $\mathcal{M}$ within the specified tolerance.
A.7.4 Whole‐body Obstacle Avoidance:
In this experiment, we control a Franka Emika robot in the presence of randomly placed spherical obstacles. The optimality distribution cost is
$\hat{c}_{\mathrm{wb}}(\hat{\bm{U}}_{t})=\sum_{h=0}^{H-1}r_{t+h}\Bigl(10\,\|\hat{\bm{x}}_{t+h}-\bm{p}_{\mathrm{target}}\|^{2}\\ +10^{6}\,\mathrm{coll\_b}(\hat{\bm{x}}_{t+h})+10^{6}\,\exp\bigl(-\tfrac{\mathrm{dist}(\hat{\bm{x}}_{t+h})}{0.01}\bigr)\\ +10\,\|\hat{\bm{u}}_{t+h}\|^{2}\Bigr)+10^{3}\,r_{t+H}\|\hat{\bm{x}}_{t+H}-\bm{p}_{\mathrm{target}}\|^{2},$ (40)
where all quantities are in joint space, $\mathrm{dist}(\cdot)$ is computed via the Robot Distance Field (Li et al., 2024), and obstacles are five spheres of radius $0.2\,\mathrm{m}$.
The feasibility expert uses a soft‐penalty distribution:
$p_{\text{feas}}\bigl(\bm{x}^{o}_{1:N_{o}},\,\bm{x}_{t},\,\bm{u}_{t}\bigr)=\prod_{\tau\in\{t,t+1\}}\Bigl(1-\exp\bigl(-\tfrac{\mathrm{dist}(\bm{x}_{\tau})}{\lambda}\bigr)\Bigr),$
where $\lambda=0.5$. This definition assigns low probability to states near obstacles, ensuring both the current and next robot configurations remain safely distant from all obstacles.
A.8 Real-robot Experiment